OpenStudy (anonymous):

(3q^2)^4 my answer 3q^8 am I right I hope ?

OpenStudy (anonymous):

When you have an exponent to a power, you multiply the exponents: \[(x ^{5})^{2}= x ^{10}\] And in your example, you're right!

OpenStudy (anonymous):

thxs

OpenStudy (campbell_st):

you need to change 3 to 3^4 abd evaluate it the power of 4 is operating on the 3 and the q

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

you have to raise everything to the power of 4, not just the \(q^2\) part

OpenStudy (campbell_st):

your answer shows the power of 4 operating on q given your answer... the question would have been 3(q^2)^4

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

believe @campbell_st he is right

OpenStudy (anonymous):

\((ab)^n=a^nb^n\) so \((3q^2)^4=3^4q^8\)

OpenStudy (anonymous):

12q^8 is the answer?

OpenStudy (anonymous):

careful here

OpenStudy (anonymous):

\(3^4\neq 3\times 4\) \[3^3=3\times 3\times 3\times 3\]

OpenStudy (anonymous):

typo there, i meant \[3^4=3\times 3\times 3\times 3\]

OpenStudy (anonymous):

oh 81 I see its 81q^8

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

wow thxs